OpenStudy (anonymous):

The secretary spent her 35 hour work week in this manner: 1/5 of her time sorting mail, 1/7 of her time filing letters, and ½ of her time typing. The rest of her time she acted as a receptionist. The percentage of her time spent as a receptionist during the week was most nearly

OpenStudy (anonymous):

no 70-59/70

OpenStudy (anonymous):

1/5 * 35 = 7 hours sorting + 1/7 * 35 = 5 hours filing + 1/2 * 35 = 17.5 hours typing -------------------------------- 29.5 hours subtotal 35 - 29.5 = 5.5 hours as receptionist 5.5 / 35 = abt. 15.7% of her time spent as receptionist.

OpenStudy (anonymous):

thank you